Translate each statement into a mathematical statement. Use k as the constant of variation:
1. M varies jointly as p and q.
2. Y varies directly as x and inversely as the square of z.
3. The time required to travel is directly proportional to the temperature and inversely proportional to the pressure P.
4. The heat H produced by an electric lamp varies jointly as the resistance R and the square of the current C.

Answer by Cromlix(4381)

Hi there,
1. M varies jointly as p and q.
M = kpq
2. Y varies directly as x and inversely as the square of z.
Y = kx/z^2
3. The time required to travel is directly proportional to the temperature and inversely proportional to the pressure P.
t = kT/P
4. The heat H produced by an electric lamp varies jointly as the resistance R and the square of the current C.
H = kRC^2
Hope this helps :-)